### v4

@@ Abstract
We report a framework that enables the broad adoption of authentic research educational methodology at various schools. We list and address common barriers that appear in many existing authentic research education programs. In our program, teams of students with complementary skills develop useful artificial intelligence (AI) solutions for researchers in natural sciences. To accomplish this, we work with research laboratories that reveal/specify their needs, and then our student teams work on the discovery, design, and development of an AI solution for unique problems using a consulting-like arrangement. To date, our group has been operating at New York University (NYU) for six consecutive semesters, has engaged more than eighty students, ranging from first-year college students to master’s candidates, and has worked with around twenty projects and collaborators. While creating education benefits for students, our approach also directly benefits scientists, who get an opportunity to evaluate the usefulness of machine learning for their specific needs.
@@ 1 Introduction
In this manuscript, we report the design and implementation of a novel AREd-type program and argue that it overcomes common AREd adoption barriers. This program was designed and implemented in the form of a group ”AI for Scientific Research” (AIfSR) at New York University, under the umbrella of Vertically Integrated Projects [^1] https://www.vip-consortium.org/ - VIP for short. Note that VIP can be used for various pedagogy approaches, and one can find reports on other programs relying on VIP machinery [12, 13]. Our focus here is not on the pedagogical outcomes of this particular program but rather on the question of how to simplify the implementation of such programs. Evaluation of AREd programs can be found elsewhere (for example, in [9, 10]).
@@ 2 AREd: definition and implementation barriers
Process of science
A student is engaged in practices built around steps and methods in traditional research (student-generated questions, hypothesis formulation, experimental design, data collection, data analysis, presentation/publication)
Novel questions
The answers to such questions are unknown and may not even exist
@@
Student-to-qualified-instructor ratio
There is a need for highly qualified (in pedagogy, material knowledge, and research methods) instructors, and a low student-to-instructor ratio
Lab/Data cost
Modern science is advanced - There is a need for advanced equipment, expensive consumables, etc.
Skills and Knowledge
Students should have sufficient skills and knowledge to tackle AREd tasks
Project formulation
Finding a good project may require significant time and expertise
Time for execution
Execution of an authentic research project takes time, which often goes beyond the time ranges available in traditional courses
@@
3 Hybrid approach to AREd
@@
As a side benefit, although an important one, the hybrid approach gives students a chance to participate in activities that can be found in real research but are usually hard to include in an AREd program if a student is supposed to focus on many aspects at once (one would need to prioritize heavily). Activities such as outreach to the ”out-of-your-domain” community, writing/journalism on scientific topics, marketing and commercialization of scientific solutions, and developing scientific code into useful ready-to-consume products are, in some sense, non-essential but do allow to bring research experience out of the purely academic silo. This is especially important given that the majority of people doing research would work outside of academia after they graduate [15]. In our framework, we demonstrate to students through practice that there are many ways to contribute to scientific research meaningfully, and each of them is important.
4 Group organization
The AI for Scientific Research (AIfSR) group is composed of bachelor’s and master’s students from NYU. The whole group is divided into small student teams (usually 3-4 students), with each of them operating as a small consulting company for a collaborating lab (Fig. 1). To address common barriers listed in the table 2, we involve:
@@ 5.1 Focus on ”AI + science”
As mentioned above, our goal is to work on ”AI + X” projects, where X is a specific area in one of the natural sciences. These areas have plenty of opportunities for relatively simple AI solutions to bring many benefits. Those scientific labs where AI/ML/advanced statistics is already widely used are not the targets for the teams.
@@ 5.3 Examples of the projects
Without going into specifics on projects’ progress and outcomes, we can list the general idea for some projects we worked with.
  • Microscopy images segmentation
    Aim: an AI algorithm able to automatically extract the shape of the cells from the optical microscope images.
  • Classification of patterns in EEG data
    Aim: an ML solution able to determine specific sleeping stages in electroencephalogram (EEG) data collected from mice.
@@
    Aim: an ML model predicting vaccination responses based on participants’ features (age, weight, etc.).
  • Focused Ion Beam (FIB) profiles
    Aim: an AI algorithm able to estimate Focused Ion Beam parameters based on the shape of the etching holes produced in a sample ([16] has more details on the topic)
6 Group Operations
@@ 7 Summary and conclusion
7.1 AREd-type program
From the information given in the manuscript, a reader can see that experience in the AIfSR group gives students an opportunity to meaningfully participate in all ”process of science” activities listed in table 1 such as student-generated questions, hypothesis formulation, experimental design, data collection, data analysis, and presentation/publication. Students participate in all these activities while working with novel questions. Thus, we believe the presented framework can be definitely classified as an AREd-type program.
@@ 7.2 Addressing barriers
In our particular case, the student-to-instructor ratio is close to 30/1. The currently required involvement of the mentor is only a few hours per week. Supervisor involvement was greater during the first semesters while the group was only starting.
@@ 7.3 Conclusion
We believe such or similar programs may be implemented at many universities. Such a program would benefit all the parties involved: students, researchers in natural sciences, mentors, as well as faculty, and administrators at universities.
8 Contributions and Acknowledgments
Sergey Samsonau leads the efforts and is responsible for the initial idea, establishment of the AIfSR group, and group development. All authors actively participated in work on projects, establishing collaborations, management, and design of group operations. Sergey Samsonau, Lu Jiang, Hazem Lashen, Jiamu Bai, Ruoxi Wang, Laiba Mehnaz, Zecheng Wang, and Ishita Patil wrote the manuscript.
We are immensely grateful to all the students who participated and are still participating in the AIfSR group, VIP program management at NYU Tandon, NYU IT, and all the researchers who collaborated with us on various projects.

### v5

@@ Abstract
We report a framework that enables the wide adoption of authentic research educational methodology at various schools by addressing common barriers. The guiding principles we present were applied to implement a program in which teams of students with complementary skills develop useful artificial intelligence (AI) solutions for researchers in natural sciences. To accomplish this, we work with research laboratories that reveal/specify their needs, and then our student teams work on the discovery, design, and development of an AI solution for unique problems using a consulting-like arrangement. To date, our group has been operating at New York University (NYU) for seven consecutive semesters, has engaged more than a hundred students, ranging from first-year college students to master’s candidates, and has worked with more than twenty projects and collaborators. While creating education benefits for students, our approach also directly benefits scientists, who get an opportunity to evaluate the usefulness of machine learning for their specific needs.
keywords:
education, authentic, research, artificial intelligence, science
@@ 1 Introduction
In this manuscript, we outline the principles of a framework that allows the design and implementation of novel AREd-type programs and argue that such a framework overcomes common AREd adoption barriers. We demonstrate the use of the principles through the discussion of the program ”AI for Scientific Research” (AIfSR) at New York University. Our focus here is not on the pedagogical outcomes of this particular program but rather on the question of how to simplify the implementation of such programs. Evaluation of AREd programs can be found elsewhere (for example, in [9, 10]).
@@ 2 AREd: definition and implementation barriers
Process of science A student is engaged in practices built around steps and methods in traditional research (student-generated questions, hypothesis formulation, experimental design, data collection, data analysis, presentation/publication)
Novel questions The answers to such questions are unknown and may not even exist
@@
Student-to-qualified-instructor ratio There is a need for highly qualified (in pedagogy, material knowledge, and research methods) instructors, and a low student-to-instructor ratio
Lab/Data cost Modern science is advanced - There is a need for advanced equipment, expensive consumables, etc.
Skills and Knowledge Students should have sufficient skills and knowledge to tackle AREd tasks
Project formulation Finding a good project may require significant time and expertise
Time for execution Execution of an authentic research project takes time, which often goes beyond the time ranges available in traditional courses
@@
3 Ten guiding principles for AREd program design
Here, we formulate the principles that we believe allow educators to design programs that are not limited by the barriers listed above.
  1. 1.
    Collaborate with researchers with expertise and data to have a basis for formulating worthy, unique projects, and who would like to explore the benefits of AI for their research
  2. 2.
    Build multidisciplinary teams of students with complementary skills and perspectives: research, AI technology, management (and others when applicable). Only accept students whose existing skills can benefit a team
  3. 3.
    Expect to use data provided by collaborators. Suggestions to collect new kinds of data may be provided, but don’t expect students to learn and use scientific equipment to produce new data
  4. 4.
    Apply best practices for work organization accepted in the industry
  5. 5.
    Treat each team of students as a small startup or a consulting team working on a proof-of-concept project: timelines, priorities, success (everyone as a unit or no-one)
  6. 6.
    Prioritize practical results/delivery (even if small) to long-term research ideas
  7. 7.
    Organize student leadership structure to enable continuity/succession of projects and collaborations, retention of expertise, promotion of idea ”we, students, have enough expertise to own and deliver”, as well as streamlined management
  8. 8.
    Use a ”Hybrid approach to AREd”: A student has exposure to and influence on all activities of the team but with a focus/priority on one of them only
  9. 9.
    Keep ownership of the project within a team, with plenty of freedom for teams to direct development. Keep an eye on and don’t allow your students to become ”free labor” for collaborator’s lab
  10. 10.
    Keep a clear focus on science and adoption of new technologies in sciences (don’t become an ”engineering shop” for various IT-related projects)
3.1 A note on ”Hybrid approach to AREd”
@@
As a side benefit, although an important one, the hybrid approach gives students a chance to participate in activities that can be found in real research but are usually hard to include in an AREd program if a student is supposed to focus on many aspects at once (one would need to prioritize heavily). Activities such as outreach to the ”out-of-your-domain” community, writing/journalism on scientific topics, marketing and commercialization of scientific solutions, and developing scientific code into useful ready-to-consume products are, in some sense, non-essential but do allow to bring research experience out of the purely academic silo. This is especially important given that the majority of people doing research would work outside of academia after they graduate [13]. As educators, we need to demonstrate to students through practice that there are many ways to contribute to scientific research meaningfully, and each of them is important.
3.2 Guiding principles in practice
Now, let us look at the details of the implementation of the AIfSR program at NYU, which was formed with these principles in mind. This will add practical details and will allow us to see why these principles accomplish the stated goal.
4 AIfSR group organization
The AI for Scientific Research (AIfSR) group was formed and is operating at NYU under the umbrella of Vertically Integrated Projects [^1] https://www.vip-consortium.org/ - VIP for short. Note that VIP can be used for various pedagogy approaches, and one can find reports on other programs relying on VIP machinery [14, 15].
The AI for Scientific Research (AIfSR) group is composed of bachelor’s and master’s students from NYU. The whole group is divided into small student teams (usually 3-4 students), with each of them operating as a small consulting company for a collaborating lab (Fig. 1). We involve:
@@ 5.1 Focus on ”AI + science”
The group’s goal is to work on ”AI + X” projects, where X is a specific area in one of the natural sciences. These areas have plenty of opportunities for relatively simple AI solutions to bring many benefits. Those scientific labs where AI/ML/advanced statistics is already widely used are not the targets for the teams.
@@ 5.3 Examples of the projects
Here is the list, which can give a general idea about some of the projects we have worked on.
Biology
@@
  • Sleep patterns classification in Electroencephalography data
  • Optical microscope images segmentation for determination of cells’ shapes
  • Image-based prediction for worms embryonic lethality
  • Classifying disease (malaria or not) based on symptoms
  • Genetics: protein sequences analysis with Large Language Model (LLM)
Chemistry
  • Determine the electric current in batteries from the magnetic field around them
  • Scoring and sampling for molecular crystal ranking
  • Cancer detection through Mass Spectrometry Imaging using ML
Physics
  • Focused Ion Beam (FIB) beam profile determination based on etching profiles
  • Analysis of motion trajectories of particles under a microscope
  • ML-enhanced analysis of spectral data collected for micro-diamonds
  • Detection and classification of pests’ presence based on acoustics data
  • AI-enabled analysis of the movement of antiferromagnetic domain walls over time
Psychology
  • Generative AI model for morphemes and connection to Magnetoencephalography data (Neurolinguistics of Morphology)
  • Use of LLMs to label generic gender categories text
Scientific Software
  • Automatic annotation generation for scientific figures
  • Addressing challenges of automatic parsing of scientific papers PDFs
  • Improving benchmarking for cloud coverage detection software with a performance evaluation on HPC (High Performance Computing) cluster
6 AIfSR group operations
@@ 7 Summary and conclusion
7.1 Is AIfSR an AREd-type program?
From the information given in the manuscript, a reader can see that experience in the AIfSR group gives students an opportunity to meaningfully participate in all ”process of science” activities listed in table 1 such as student-generated questions, hypothesis formulation, experimental design, data collection, data analysis, and presentation/publication. Students participate in all these activities while working with novel questions.
Thus, we believe the presented framework principles allowed us to design an AREd-type program.
@@ 7.2 Addressing barriers
In our particular case, the student-to-instructor ratio is close to 30/1. The currently required involvement of the mentor is several hours per week. Supervisor involvement was greater during the first semesters while the group was only starting.
@@ 7.3 Conclusion
We believe that the principles outlined in this manuscript will allow similar programs to be implemented at many universities. Such programs would benefit all the parties involved: students, researchers in natural sciences, mentors, as well as faculty, and administrators at universities.